# EEG Emotion Copilot: Optimizing Lightweight LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

### v2

@@ V-A Data Structure of Prompt
We developed a comprehensive human-based dataset that integrates both public knowledge datasets and specialized emotion datasets, facilitating the construction of a robust data structure to drive the development of LLMs. By employing effective prompting strategies, we enhanced the model’s ability to generate contextually relevant responses. Furthermore, while the performance of the models may not be exceptionally high, as illustrated in Table I and II, language remains more challenging to cluster unsupervised compared to other data types [63]. Nonetheless, by utilizing a pretrained sentence transformer model of an appropriate size, the model is still able to capture the underlying emotional features of the language to some extent as demonstrated in the spectral clustering shown in Fig. 9, despite potential issues with feature imbalance.
V-B Data Redundancy in EEG Signal Processing
We addressed the issue of data redundancy inherent in EEG signal processing. Through efficient data compression techniques, we optimized the handling of long sequences of EEG data, significantly improving computational efficiency and facilitating real-time emotion computation. However, simply compressing signal data is computationally intensive in many cases, and in complex scenarios, more channel signals are needed. Therefore, if LLMs can generate dense channel signals using only a few channel data and then analyze the emotions of the subjects [64], [65], computational efficiency will be greatly improved.
V-C Patient Privacy and Model Efficiency
@@
V-D Performance of Lightweight Models
@@
We performed model pruning and compared the performance of the models obtained through the two fine-tuning methods illustrated in Fig. 5. It was found that the results obtained using Strategy 1 for fine-tuning were superior. Furthermore, as shown in Table I and II, the lightweight pruned model achieved through multi-step fine-tuning demonstrated better performance compared to other models with larger parameter sizes.
V-E Hallucination of Heavyweight Models
@@
While a low training loss suggests a good fit to the training data, it does not necessarily reflect the model’s true performance, as the generated responses may still fail to meet expectations or even diverge significantly [66]. This issue is particularly evident in 7B and larger models, where the hallucination problem becomes more pronounced in specialized domains. These findings also suggest that lightweight models may be better suited for such domain-specific generation tasks.
VI Conclusion
@@
In this paper, we present EEG Emotion Copilot, a lightweight and locally-running LLM designed for emotional recognition using EEG signals and the generation of corresponding treatment plans for assisted electronic medical records. Our findings indicate that EEG Emotion Copilot has the potential to revolutionize the recognition and treatment of emotions in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will combine lightweight LLMs with multimodal data such as human facial expressions, speech, gestures to achieve greater accuracy in affective computing.

### v3

@@ 5.1 Data Structure of Prompt
We developed a comprehensive human-centered dataset that integrates both public knowledge bases and specialized emotion datasets, establishing a robust data foundation for LLM development. By leveraging effective prompting strategies, we enhanced the model’s ability to generate contextually relevant responses. While the overall model performance may not be exceptionally high, as shown in Tables 2 and 3, unsupervised clustering of language remains inherently more challenging than other data types [76]. Nevertheless, by employing a pretrained sentence transformer of an appropriate size, the model can still capture underlying emotional features to some extent, as demonstrated by the spectral clustering results in Fig. 11, despite potential feature imbalance.
5.2 Data Redundancy in EEG Signal Processing
To address the issue of data redundancy inherent in EEG signal processing, we employ wavelet-based compression techniques to optimize the handling of long EEG sequences, significantly improving computational efficiency and enabling rapid emotion recognition. However, direct signal compression remains computationally intensive in many scenarios, particularly when complex affective states require information from more EEG channels. To this end, recent advances suggest that large language models (LLMs) can be leveraged to generate dense, high-fidelity EEG signals from a limited number of input channels, thereby enabling efficient downstream emotion analysis [77], [78]. Furthermore, by integrating channel-aware prompts, this approach mitigates inter-subject and temporal variability in EEG data, ensuring robust and generalizable emotion recognition across diverse clinical contexts.
5.3 Patient Privacy and Model Efficiency
@@
5.4 Performance of Lightweight Models
@@
We performed model pruning and compared the performance of the models obtained through the two fine-tuning methods illustrated in Fig. 5. It was found that the results obtained using Strategy 1 for fine-tuning were superior. Furthermore, as shown in Table 2 and 3, the lightweight pruned model achieved through multi-step fine-tuning demonstrated better performance compared to other models with larger parameter sizes.
Table 6: Comparison of inference time and memory usage among different models. Inference time is reported for both Top-$k=50$ and Top-$k=1$ decoding.
Model Inference Time Memory Used
Top-$k=50$ Top-$k=1$ (MB)
ours 26.002s 60.629s 308.912
qwen2_0.5B 51.362s 132.187s 959.074
opt-350m 123.165s 173.729s 647.707
LiteLlama 74.044s 74.533s 888.567
gpt-sw3-356m 128.207s 186.361s 813.613
qwen3-0.6B 125s 437.262s 1156.125
qwen2-1.5B 155.559s 123.165s 2979.620
Internlm2.5-1.8B 434.379s 420.147s 3633.197
qwen2.5-3B 317.598s 258.132s 5943.055
DS-R1-Dis-Qwen-1.5B 322.975s 294.173s 3424.745
qwen2.5-7B 1004.48s 3283.745s 14602.636
Internlm2.5-7B 16475s 5143.517s 14830.516
Figure 16: BLEU of Various Models on MHCCS.
5.5 Hallucination of Heavyweight Models
@@
While a low training loss suggests a good fit to the training data, it does not necessarily reflect the model’s true performance, as the generated responses may still fail to meet expectations or even diverge significantly [79]. This issue is particularly evident in 7B and larger models, where the hallucination problem becomes more pronounced in specialized domains. Heavyweight models, despite their extensive pre-training on general datasets like Wikitext [26] and Common Crawl [27], often struggle to adapt to the nuanced patterns of compressed EEG signals and the integration of demographic data required for accurate emotion classification and electronic medical record (EMR) generation. This is evident in their lower F1 scores for nine-emotion recognition (e.g., 0.169–0.179 for Qwen2.5-7B and Internlm2.5-7B, Table 3), which are only marginally above the random-guess baseline of 0.111. The hallucination problem is exacerbated by the models’ tendency to overfit to general language patterns, leading to outputs that are less precise in specialized contexts. For example, qualitative analysis of EMRs generated by Qwen2.5-7B revealed occasional inclusion of irrelevant treatment suggestions, such as generic recommendations not aligned with the inferred emotional state, whereas our model produced more concise and task-relevant outputs. These findings also suggest that lightweight models may be better suited for such domain-specific generation tasks.
5.6 Future Work
Future work will focus on advancing the system’s clinical applicability and accuracy. First, we plan to integrate lightweight LLMs with multimodal data, such as facial expressions, speech, and gestures, to improve the robustness and effectiveness of affective computing. Second, to streamline clinical workflows, we will develop a module to interface with hospital EMR databases or patient ID systems, automatically retrieving demographic information (e.g., gender, age) to eliminate redundant user input. Finally, to ensure clinical trustworthiness, we will implement a human-in-the-loop interface for physicians to review and refine diagnostic suggestions, incorporating patient symptoms and clinical expertise, and conduct a clinical validation study to evaluate diagnostic accuracy in real-world settings. These enhancements will solidify the EEG Emotion Copilot’s role as a reliable assistive tool in medical practice.
6 Conclusion
@@
In this paper, we introduce EEG Emotion Copilot, a light weight, locally-executable LLM designed for emotion recognition from EEG signals and the generation of corresponding treatment plans to assist in electronic medical record management. Our findings suggest that EEG Emotion Copilot holds transformative potential in advancing emotion recognition and treatment in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will explore the integration of lightweight LLMs with multimodal data, including facial expressions, speech, and gestures, to further elevate the accuracy and effectiveness of affective computing.
CRediT authorship contribution statement
Hongyu Chen: Writing–original draft, Methodology, Formal analysis, Conceptualization, Software. Weiming Zeng: Writing–review & editing, Supervision, Project administration, Funding acquisition. Chengcheng Chen: Writing–review & editing. Luhui Cai: Writing–review & editing. Fei Wang: Writing–review & editing. Yuhu Shi: Writing–review & editing. Lei Wang: Writing–review & editing. Wei Zhang: Writing–review & editing. Yueyang Li: Writing–review & editing. Hongjie Yan: Writing–review & editing. Wai Ting Siok: Writing–review & editing. Nizhuan Wang: Writing–review & editing, Supervision, Project administration, Funding acquisition.
Declaration of competing interest
The authors declare that they have no competing interests.
Data availability
The datasets are publicly available.
Code availability
The Code will be released at
https://github.com/NZWANG/EEG_Emotion_Copilot.
Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund [Project ID: P0053210], The Hong Kong Polytechnic University Faculty Reserve Fund [Project ID: P0053738], an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428), The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097) and the National Natural Science Foundation of China [grant number: 31870979].